OpenStudy (anonymous):

how can i get the answer 24 by using the numbers 8,8,3,3

OpenStudy (lgbasallote):

what operation are you allowed to use?

OpenStudy (saifoo.khan):

im getting 24.333

OpenStudy (lgbasallote):

you can do\[\large (\sqrt{8 \times 8})(\sqrt{3 \times 3})\]

mathslover (mathslover):

is it multiplication ?

OpenStudy (lgbasallote):

i think it's a riddle...

OpenStudy (lgbasallote):

well what i said is all i can think of @shaziazafar =))

mathslover (mathslover):

if it is :square root of multiplication of the numbers then : \[\large{\sqrt{8\times 8\times 3 \times 3}}\] \[\large{\sqrt{8^2\times 3^2}}\] \[\large{\sqrt{8^2}\times \sqrt{3^2}}\] \[\large{8^{{\frac{1}{2}}*2}\times 3^{{\frac{1}{2}}*2}}\] \[\large{8^{\frac{\cancel{2}^1}{\cancel{2}^1}} *3^{\frac{\cancel{2}^1}{\cancel{2}^1}}}\] \[\large{8*3=24}\]

mathslover (mathslover):

"Operation matters a lot"

OpenStudy (saifoo.khan):

Didn't @lgbasallote had same?

mathslover (mathslover):

@lgbasallote as per the question of @shaziazafar he asked for the operation..

OpenStudy (lgbasallote):

what i did was operation =_=

OpenStudy (lgbasallote):

and there was also no specification of operation

mathslover (mathslover):

So the answer is : square root of multiplication of the numbers @saifoo.khan I represented the whole way thanks!!

mathslover (mathslover):

I agree i think that the asker must have explained what he wanted as the answer..

OpenStudy (lgbasallote):

that's what made his answer better

OpenStudy (lgbasallote):

maybe if i do \[(\sqrt{8 \times 3}) \times(\sqrt{3 \times 8} )\implies \sqrt {24} \times \sqrt{ 24} \implies \sqrt{24 \times 24} \implies \sqrt{(24)^2} \implies 24\] i'll have a good answer too :o

mathslover (mathslover):

:D ... You have best response dude:)

OpenStudy (lgbasallote):

"good answer" literally lol..not the medal

mathslover (mathslover):

:D .. sorry for misunderstanding

OpenStudy (lgbasallote):

actually... 8 + 8 + 3 + 3 = 22

OpenStudy (ganpat):

oops...

OpenStudy (unklerhaukus):

\[3\times8=24\]